OpenStudy (anonymous):

An element X has two isotopes. X-15 has 7 protons and 8 neutrons. Determine the atomic number and mass number of the second isotope, which has 10 neutrons. A) atomic number=8; mass number=9 B) atomic number=8; mass number=17 C) atomic number=7; mass number=17 D) atomic number=7; mass number=15

OpenStudy (anonymous):

I don't understand how to get the answer, can anyone help me.....

OpenStudy (anonymous):

Yes brother, sure!

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Just to start with, I would like to point out to you that isotopes are elements of the same compound that have the same proton number, but different neutron and consequently different mass numbers. According to your problem, if the second isotope has 10 neutrons, still has to have 7 protons. That is, in chemistry's periodic table, an element is identified by the particular proton number. And as I said before, isotopes are of the same element, the reason why their proton number is the same, to keep the identity, regardless the electron or neutron number. Answer would be C.

OpenStudy (anonymous):

Oh so I need a Periodic table to solve it???

OpenStudy (anonymous):

For this exercise no, because you should know that isotopes have the same proton number, regardless of the change in neutron or electron number. In regard of periodic table was my explanation.

OpenStudy (anonymous):

I'm not understanding is there like a mathematical equation that I have to do to get it?

OpenStudy (anonymous):

Mass number = Proton/atomic number + Neutron number. Does that explain?

OpenStudy (anonymous):

No I'm still not getting the answer: protons/ atomic number+ neutrons I'm not understanding this at all I'm sorry 😓

OpenStudy (anonymous):

To find the mass number of an element, you have to sum the number of protons with the number of neutrons. For instance, using the element X given in the statems, the mass number is 15, because the proton number (7) + the neutron number (8) equals the mass 15. How is this one?

OpenStudy (anonymous):

So 15 is the mass number because 7+8=15?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Okay so then the answer is D?

OpenStudy (anonymous):

No, C. Remember they told you the second isotope has a neutron number 10. So it becomes 7+10=17.

OpenStudy (anonymous):

Okay so you take the 8 out?

OpenStudy (anonymous):

No. Given that the proton number of isotopes is the same, the neutron number is what changed, hence 7+10=17.

OpenStudy (anonymous):

Okay so what do you do with the 8 then?

OpenStudy (anonymous):

Sorry it took so long to reply back I didn't have wifi

OpenStudy (photon336):

Mass number = protons + neutrons Z Atomic number = number of protons A An element X has two isotopes. X-15 has 7 protons and 8 neutrons. Determine the atomic number and mass number of the second isotope, which has 10 neutrons. the number of protons = identity of your element which will not change, you change the number of protons, you'll get a different element. so number of protons stays the same. your asked for the atomic number and mass number for the second isotope based on my first argument you should have 7 protons as the atomic number then if something has 10 neutrons you sum the numbers 7+10 = 17 So for an isotope, which is just the same element just with a different number of neutrons.

OpenStudy (photon336):

most important is the fact that isotopes, are of the same element so their number of protons will be the same but will differ in # of neutrons.

OpenStudy (anonymous):

Thank you both